Determine the missing amounts. Enter decreases as negative amounts.

Apple Verizon
Year 1:
Assets $231,839 $
Liabilities $ $
Stockholders' equity $ $ 38,836
Increase (Decrease) in assets, liabilities,
and stockholders' equity during Year 2:
Assets $ 58,640 $
Liabilities 50,832 $(14,852)
Stockholders' equity 7,808 $
Year 2:
Assets $ $232,708
Liabilities $171,124 $
Stockholders' equity $ 12,298
Answer #1

Apple

Year 1

Assets $231,839

Liabilities =(171124-50832)=$120292

Stockholders' equity=Assets -Liabilities =$231,839-$120292=$111547

Year 2:

Assets =2321839+58640=$290479

Liabilities=$171124

Stockholders' equity=$290479-171124=$119355

Verizon

We have to start form year 2

Year 2:

Liabilities=Assets -Stockholders' equity =232,708-12298=$220410

Stockholders' equity =12,298

Year 1:

Liabilities =220410-14852=$205558

Assets=Liabilities+Stockholders' equity =38836+$205558=$244394

Increase (Decrease) in assets, liabilities,
and stockholders' equity during Year 2:

Assets=$244394-232708=$-11686

Stockholders' equity=$38836-12298=$26358

Increase (Decrease) in assets, liabilities,
and stockholders' equity during Year 2: